# Drawback of Enforcing Equivariance and its Compensation via the Lens of Expressive Power

### v1

@@ Abstract
Equivariant neural networks encode symmetry as an inductive bias and have achieved strong empirical performance in wide domains. However, their expressive power remains not well understood. Focusing on 2-layer ReLU networks, this paper investigates the impact of equivariance constraints on the expressivity of equivariant and layer-wise equivariant networks. By examining the boundary hyperplanes and the channel vectors of ReLU networks, we construct an example showing that equivariance constraints could strictly limit expressive power. However, we demonstrate that this drawback can be compensated via enlarging the model size. Furthermore, we show that despite a larger model size, the resulting architecture could still correspond to a hypothesis space with lower complexity, implying superior generalizability for equivariant networks.
@@ 1 Introduction
In scaling up the capacity of machine learning models, leveraging the intrinsic symmetry in data, if it exists, would be a much more efficient approach than simply scaling up model size. Good examples of this intrinsic symmetry would be point cloud processing qi2017pointnet; li2018pointcnn; fuchs2020se; chen2021equivariant, graph neural networks velivckovic2017graph; kanatsoulis2024graph, and chemistry (faber2016machine; eismann2021hierarchical). When learning these symmetries, vanilla neural network-based methods must be shown numerous (sometimes manually) transformed data examples, such as rotated or translated images, and enforced to make invariant/equivariant predictions, which is usually computationally expensive and requires vast amounts of data (bronstein2021geometric). Equivariant neural networks overcome these challenges by constraining their output to be invariant or equivariant to group representations (cohen2016group), through either new architectures or optimization methods. One useful approach is to enforce each layer of the network to be equivariant.
While the equivariance constraint in a neural network guarantees the symmetry properties, it remains unclear whether they compromise the expressive power. In this paper, we address this by determining whether equivariance (or layer-wise equivariance) increases the expected risk when approximating an equivariant target function. Specifically, we examine if the following strict inequality holds for some equivariant target function $s$ and some parameter spaces $\Theta$:
@@
where $\Theta$ is the set of network parameters (usually $L$ matrices) and $F_{\theta}$ is the neural network with parameter $\theta$. Notation GENs means the model parameters that make $F_{\theta}$ an equivariant network, while LENs is the parameter set of layer-wise equivariant networks.
@@
Interpretation of Equivariant Constraints.
The equivariance constraints primarily govern the input-output mapping. To understand how this affects the model parameters, we focus on the required weight matrices. We demonstrate that for 2-layer ReLU networks, requiring equivariance forces the network to have specific structural symmetries. For GENs, this necessity manifests as the boundary hyperplanes that define the network’s linear regions must be symmetric. While the layer-wise equivariance constraint (LENs) requires sufficient model capacity to contain symmetric channel vectors. This framework, which connects the functional requirement of equivariance to its necessary structural consequences, provides a clear basis for comparing their expressive power.
Drawback of layer-wise equivariance in expressive power.
Following the insight of how equivariance restricts model parameters, we demonstrate a scenario where equivariance strictly hurts the expressive power. Specifically, we construct an invariant target function $s$ on $\mathbb{R}^{2}$ and establish two key findings: 1) for a 1-neuron 2-layer architecture, networks without equivariance constraints achieve a strictly lower expected error compared to GENs, and 2) for a 2-layer architecture, LENs require a strictly larger model size than GENs to fully approximate the target function.
@@
Inversely, we prove that an increase in model size can fix the reduced expressive power caused by (layer-wise) equivariance. When the target function is invariant, and the data distribution is symmetric, enlarging the model size of any equivariant network by $|G|$ times can compensate for the expected loss over a 2-layer neural network. Moreover, a doubled model size allows LENs to represent all invariant functions of GENs. We further construct a LEN with $|G|$ times enlarged model size whose hypothesis space has a lower dimension. This implies that the LEN has improved generalizability; together with the comparable expressive power, this result renders LENs a competitive solution.
To the best of our knowledge, this paper provides the first in-depth theoretical examination of the expressive power of layer-wise equivariant neural networks. We also propose that an appropriately larger model size can compensate for the loss in expressive power and potentially enhance model generalisability, underlining the advantages of incorporating layer-wise equivariance into model architectures.
@@ 2 Related Works
Designing equivariant neural networks. Remarkable advances have been seen in designing equivariant network architectures for the symmetries in data. Typical approaches for obtaining an equivariant neural network can be designing equivariant architectures cohen2016steerable; satorras2021n or introducing equivariance constraints into the optimization winter2022unsupervised; tang2022equivariance.
@@
Equivariant neural architectures. A typical designing paradigm is designing equivariant neural architectures based on classic neural networks. For example, cohen2016group design group convolutional neural networks in the group spaces as the equivariant variants of the convolutional neural networks shin2016deep, which is extended to the homogeneous spaces by cohen2019general. Similar approaches are also seen in designing equivariant graph neural networks klicpera2020directional, equivariant transformers hutchinson2021lietransformer, equivariant diffusion hoogeboom2022equivariant, equivariant point network chen2021equivariant, equivariant harmonic networks worrall2017harmonic, etc.
@@
Symmetry-aware optimization. In this stream, equivariance constraints on the the parameters of the networks are introduced in the optimization to ensure the equivariance of the learned model, usually called steerable neural networks cohen2016steerable. For examples, methods surge to compute the solution space satisfying the equivariance constraints for different groups and network architectures, including the symmetry group $S_{N}$ in steerable graph networks maron2018invariant and the rotation and Euclidean groups $\{C_{N},E_{2}\}$ in steerable convolutional neural networks weiler2018learning; weiler2019general. A major obstacle is the high computational costs to compute the equivariant basis. finzi2021practical address this issues with an algorithm with polynomial computational complexity on the sum of the number of discrete generators and the dimension of the group, which divides the problem into some independent subproblems and adopts Krylov method to compute nullspaces.
@@
Theoretical analyses. With many works contributing in deep learning theory he2020recent, theoretical studies have shown that equivariant neural networks have significant advantages from the aspects of generalizability, convergence, and approximation. sannai2021improved prove improved generalization error bounds of equivariant models . lawrence2021implicit prove that linear group convolutional neural networks trained by gradient descent for binary classification converge to solutions with low-rank Fourier matrix coefficients based on the results via implicit bias. zaheer2017deep; maron2019universality; yarotsky2021universal prove the universal approximation ability of equivariant models by various approximation techniques, while ravanbakhsh2020universal utilize group averaging to obtain an equivariant approximator to prove the universal approximation ability. In addition, elesedy2021provably prove that invariant/equivariant models have a smaller expected loss when the target function is invariant/equivariant. azizian2021expressive study the expressive power of invariant/equivariant graph neural networks.
@@ 3 Problem Settings
In this section, we introduce our problem setting. We focus on two-layer ReLU networks with invariant output functions to show how equivalence and layer-wise equivalence impact the expressive power. Besides, we also assume the groups to be finite.
General networks (GNs)
GNs are neural networks without other constraints. In this paper, we focus on the architecture of a two-layer network, which forms a function $F:\mathbb{R}^{n}\to\mathbb{R}^{d}$. The network consists of a hidden layer with weight matrix $W^{(1)}\in\mathbb{R}^{n\times m}$ and an output layer with weight matrix $W^{(2)}\in\mathbb{R}^{m\times d}$ where $m$ is the size of the hidden layer and the number of neurons. With ReLU nonlinearity $\sigma$, the input $x\in\mathbb{R}^{n}$ is mapped to the output space by $F(x)=W^{(2)}\sigma(W^{(1)}x)$. For simplicity, we represent $\alpha_{i}^{\top}$ as the $i$-th column of $W^{(1)}$ and $\beta_{j}$ as the $j$-th row of $W^{(2)}$ and $W^{(1)}$. Then the output of a GN can be formulated as follows:
@@
General equivariant networks (GENs)
GENs represent a subset of GNs that constrain the output function $F$ to be equivalent with respect to a given group representation. A group representation for a group $G$ is formally defined as a homomorphism $\rho:G\to GL(m)$, where $GL(m)$ is the group of all invertible matrices on $\mathbb{R}^{m\times m}$. A function $F$ is considered equivariant if, for given input and output representations $\rho$ and $\phi$ respectively, it adheres to the condition $F\circ\rho_{g}=\phi_{g}\circ F$ for all group elements $g\in G$. A GN whose function $F$ satisfies this property is categorised as a GEN. Specifically, when $\phi=\text{id}$, the function $F$ is called an invariant function and the networks are called invariant networks. We note that GENs aim to represent all possible equivariant networks, and assume nothing but the equivariant output functions only.
Layer-wise equivariant networks (LENs)
LENs represent a subset of GNs that constrain each layer of the network to be equivalent with respect to a given group representation. Specifically, given the group representation $\psi$ in the hidden layer, LENs require that for all $g\in G$
@@
While each layer is constrained to be equivariant to maintain the symmetry in the data, the whole network is therefore equivariant. In fact, the output function is ensured to be equivariant under the group representations $\rho$ and $\phi$, because for every group element $g\in G$, we have
$F\circ\rho_{g}$ $=W^{(2)}\circ\sigma\circ W^{(1)}\circ\rho_{g}=W^{(2)}\circ\sigma\circ\psi_{g}\circ W^{(1)}$
$=W^{(2)}\circ\psi_{g}\circ\sigma\circ W^{(1)}=\phi_{g}\circ W^{(2)}\circ\sigma\circ W^{(1)}=\phi_{g}\circ F$
In practice, constructing LENs is a common approach to achieving GENs (he2021efficient). Different from GENs, LENs take more constraints. Firstly, one group representation $\psi_{g}$ making $\sigma\circ\psi_{g}=\psi_{g}\circ\sigma$ is needed for all $g\in G$, where we call $\psi$ admitted. Besides, both $W^{(1)}$ and $W^{(2)}$ should be equivalent under some representations, where we call a matrix in the intertwiner space. Given the group representations $\rho$ and $\phi$, to construct LENs is to find an admitted group representation $\psi$ and the corresponding intertwiner space.
Expressive power
In this paper, we focus on the expressive power of GNs, GENs, and LENs, which is defined as the minimum expected loss for approximating a target function $s$, and a smaller expected loss means a better expressive power. Since GENs serve as a subset of GNs and LENs form more constraints, trivially, we have for any target function $s$, any input distribution $P\in\Delta\mathbb{R}^{n}$, and any number of neurons $m$:
@@
where $\theta$ represents the parameter of a two-layer ReLU network ($\theta=(W^{(2)},W^{(1)})$), and $F_{\theta}$ is the forward function given parameter $\theta$, and $\Theta_{m}$ is the parameter set of $m$-hidden-layer networks. We reuse the notations GNs (also for GENs and LENs) to denote the respective parameter sets that makes $F_{\theta}$ a GN (also for GEN and LEN). To question whether equivariance limits the expressive power of two-layer network is to question whether there exists some equivariant function $s$ and $m$ such that
$\inf_{\theta\in\Theta_{m}\cap\text{GNs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]$ $<\inf_{\theta\in\Theta_{m}\cap\text{GENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]$
$\text{and}~~~~\inf_{\theta\in\Theta_{m}\cap\text{GENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]$ $<\inf_{\theta\in\Theta_{m}\cap\text{LENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}].$
For the first part, it questions whether equivariance is necessary for being the optimal approximator if the target function is equivariant. As for the second part, the question becomes whether layer-wise equivariance is the optimal way of incorporating equivariance. We also note that it is reasonable to assume the target function $s$ to be equivariant, because equivariant networks are designed for equivariant tasks.
4 Symmetric Boundary Hyperplanes and Symmetric Channel Vectors
To establish a clear foundation for our analysis of expressive power, we define two new terms, symmetry boundary hyperplanes and symmetry channel vectors.
4.1 GENs implies Symmetric Boundary Hyperplanes
Since the ReLU non-linearity is a piece-wise linear function, neural networks with only ReLU as activity function are also a piece-wise linear function. Specifically, we have the following formula for two-layer networks:
@@ 4.1 GENs implies Symmetric Boundary Hyperplanes
where $\mathcal{F}(x)=\sum_{i:\left\langle\alpha_{i},x\right\rangle\geq 0}^{m}\alpha_{i}\beta_{i}^{\top}$ is called the feature function of $F$ and is piece-wise constant.
@@
It is worth noting that we have no necessary to define the feature function $\mathcal{F}$ on the boundary hyperplanes. For any given function $F$, we can define it only on the set $S=\{x:F\text{ is linear in }B(x,r),\exists r>0\}$, where $B(x,r)=\{x+z:\|z\|<r\}$ is the open ball centering at $x$ with radius $r$. The value of $\mathcal{F}(x)$ is determined uniquely by the Riesz representation theorem. The complementary set $S^{c}$ is the union of all boundary hyperplanes, whose intersection with the hyperplanes $M\pm\delta v$.
For all $x\in S$ that satisfies $\langle x,\alpha_{i}\rangle\not=0$ for all $i\in[m]$, $\mathcal{F}(x)$ must be $\sum_{i:\langle x,\alpha_{i}\rangle>0}\beta_{i}\alpha_{i}^{T}$. Therefore, if the feature function changes when crossing through same hyperplane, the active set $\{i:\langle x,\alpha_{i}\rangle>0\}$ must change and the increment of the feature function is $\sum_{i:\langle\alpha_{i},x\rangle=0}\beta_{i}\alpha_{i}^{T}$, which is then summarized as Theorem 4.2, which can be viewed as an equivalent definition by the channel vectors.
@@
Boundary hyperplane provides a powerful analytical tool for understanding the internal mechanisms of GENs. We prove that for a GEN, the set of its boundary hyperplanes are symmetric with respect to all group representations. Specifically, for any given boundary hyperplane $M$ and all fixed group representation $\rho$, applying the transformation $\rho$ to the boundary hyperplane (denoted as $\rho M=\{\rho x:x\in M\}$) also gets a boundary hyperplane, because:
@@
4.2 LENs further implies Symmetric Channel Vectors
Compared with GENs, LENs enforce the symmetry constraints on every layer of the networks. While the hyperplane symmetry previously discussed is a property of all GENs, these further constraints imposed by the LEN structure lead to additional insights. Specifically, as defined before, LENs require one admitted group representation $\psi$ for ReLU nonlinearity and two matrices in the corresponding intertwining space.
We firstly consider the property of admitted group representation, i.e., $\psi\circ\sigma=\sigma\circ\psi$. This commutation property is not trivial and places a significant restriction on the form that the group representation $\psi$ can take. Specifically, for $\sigma=\operatorname{ReLU}$, the following lemma provides a complete characterization of all such admitted representations.

### v2

@@ Abstract
Equivariant neural networks encode the intrinsic symmetry of data as an inductive bias, which has achieved impressive performance in wide domains. However, the understanding to their expressive power remains premature. Focusing on 2-layer ReLU networks, this paper investigates the impact of enforcing equivariance constraints on the expressive power. By examining the boundary hyperplanes and the channel vectors, we constructively demonstrate that enforcing equivariance constraints could undermine the expressive power. Naturally, this drawback can be compensated for by enlarging the model size – we further prove upper bounds on the required enlargement for compensation. Surprisingly, we show that the enlarged neural architectures have reduced hypothesis space dimensionality, implying even better generalizability.
@@ 1 Introduction
In scaling up the capacity of machine learning models, leveraging the intrinsic symmetry in data, if it exists, would be a much more efficient approach than simply scaling up model size; see good examples in point cloud processing (qi2017pointnet; li2018pointcnn; fuchs2020se; chen2021equivariant), visual tasks (cohen2016group; xu20232), graph neural networks (velivckovic2017graph; kanatsoulis2024graph), physics and chemistry (faber2016machine; eismann2021hierarchical; kondor2025principles), reinforcement learning and decision making (wangmathrm; qin2022benefits), amongst others. When learning these symmetries, classic methods must be shown numerous transformed (sometimes manually) data examples, such as rotated or translated images, and be enforced to make invariant/equivariant predictions, which is usually computationally expensive and requires vast amounts of data (bronstein2021geometric). Equivariant neural networks address these challenges through either new architectures or optimization methods to enforce symmetries (cohen2016group), in which a useful approach is to enforce each layer of the network to be equivariant.
Despite the impressive performance, it remains unclear whether enforcing equivariance comes at the cost of reduced expressive power. This paper investigates this trade-off by analyzing whether restricting the hypothesis space to equivariant models increases the expected risk when approximating an equivariant target function $s(x)$. Specifically, we examine whether the following inequalities hold:
@@
Here, $\Theta$ is the set of network parameters (usually $L$ matrices) and $F_{\theta}$ is the neural network with parameter $\theta$. We consider two types of equivariant networks: GENs (General Equivariant Networks), where the global mapping $x\to F_{\theta}(x)$ is required to be equivariant, and LENs (Layer-wise Equivariant Networks), a stricter subset where equivariance is enforced at every layer of the network. In practice, constructing LENs is a common approach to achieving GENs (he2021efficient).
@@
Boundary hyperplanes and channel vectors.
We first investigate how these equivariance constraints (imposed at the level of input-output mapping) fundamentally restrict the architectural structure of neural networks. For two-layer ReLU networks, we focus on the geometric properties of the boundary hyperplanes and channel vectors of equivariant networks. We prove that GEN requires boundary hyperplanes to be symmetric. While for LENs, we demonstrate a constructive property: any LEN can be reformulated into another LEN that maintains the same model size but possesses a symmetric set of channel vectors. These geometric rigidities serve as the foundation for our subsequent analysis, enabling us to quantify how much model size is enough for approximating a target function.
Drawback of enforcing equivariance.
Following the insight of the above analysis, we demonstrate a scenario where equivariance strictly hurts the expressive power. Specifically, we construct an invariant target function $s$ on $\mathbb{R}^{2}$ and establish two key findings: (1) for two-layer ReLU networks with a single neuron, general networks (GNs) achieve a strictly lower expected error compared to general equivariant networks (GENs), and (2) for two-layer ReLU networks that can arbitrarily approximate the target function, layer-wise equivariant networks (LENs) require a strictly larger model size than GENs.
@@
Inversely, we prove that an increase in model size can fix the compromised expressive power. When the target function is invariant, and the data distribution is symmetric, both in terms of group $G$, enlarging the model size of any equivariant network by $|G|$ times can compensate for the expected loss over a 2-layer neural network. Moreover, a doubled model size allows LENs to represent all invariant functions of GENs. We further prove that the constructed LEN with $|G|$ times enlarged model size surprisingly has a less complex hypothesis space, implying even improved generalizability. Together with the comparable expressive power, this result renders LENs a competitive solution.
To the best of our knowledge, this paper provides the first in-depth theoretical examination of the expressive power of layer-wise equivariant neural networks, demonstrating that enforcing equivariance has drawback in expressive power. We also show that an appropriately larger model size can compensate for the loss in expressive power and potentially enhance model generalisability, underlining the advantages of incorporating layer-wise equivariance into model architectures.
@@ 2 Related Works
Typical approaches for obtaining an equivariant neural network are (1) designing equivariant architectures (cohen2016steerable; satorras2021n; trang20243) and (2) introducing equivariance constraints into the optimization (winter2022unsupervised; tang2022equivariance; pertigkiozoglou2024improving).
Equivariant neural architectures.
@@
cohen2016group; shin2016deep design group convolutional neural networks in the group spaces as the equivariant variants of the convolutional neural networks, which is extended to the homogeneous spaces by cohen2019general. Similar approaches have also been employed in the design of other popular equivariant networks, such as graph neural networks (klicpera2020directional; aykent2025gotennet), transformers (hutchinson2021lietransformer; islam2025platonic), and diffusion models (hoogeboom2022equivariant; wan2025equivariant).
Symmetry-aware optimization.
@@
In this stream, equivariance constraints are introduced in optimization, as either regularisers or constraints, to ensure the equivariance, usually termed as steerable neural networks (cohen2016steerable). Other good examples are about enforcing the symmetry group $S_{N}$ in steerable graph networks (maron2018invariant), the rotation and Euclidean groups $\{C_{N},E_{2}\}$ in steerable convolutional neural networks (weiler2018learning; weiler2019general), and Lie symmetry in differential equations (akhound2023lie; jianglie). A major obstacle is the high computational costs to compute the equivariant basis. finzi2021practical address this issues with an algorithm with polynomial computational complexity on the sum of the number of discrete generators and the dimension of the group, which divides the problem into some independent subproblems and adopts Krylov method to compute nullspaces.
Theoretical analysis.
@@
Theoretical studies have shown that equivariant neural networks have significant advantages, in terms of generalizability, convergence, and approximation. sannai2021improved prove improved generalization error bounds of equivariant models. qin2022benefits leverage orbit averaging to construct a projection mapping from the original hypothesis space to the permutation-equivalent counterpart, underpinning reduced hypothesis complexity and better generalizability. lawrence2021implicit prove that linear group convolutional neural networks trained by gradient descent for binary classification converge to solutions with low-rank Fourier matrix coefficients based on the results via implicit bias. zaheer2017deep; maron2019universality; yarotsky2021universal; pacini2025universality prove the universal approximation ability of equivariant models by various approximation techniques, while ravanbakhsh2020universal utilize group averaging to obtain an equivariant approximator to prove the universal approximation ability. In addition, elesedy2021provably prove that invariant/equivariant models have a smaller expected loss when the target function is invariant/equivariant. azizian2021expressive study the expressive power of invariant/equivariant graph neural networks.
@@ 3 Problem Settings
General networks (GNs).
GNs refer to general neural networks without any design for realizing the equivariance. In this paper, we focus on two-layer networks, with a hidden layer of width $m$ (also the number of neurons), representing a mapping from $\mathbb{R}^{n}\to\mathbb{R}^{d}$. The network is parameterized by $\theta$, including the weight matrix $W^{(1)}\in\mathbb{R}^{n\times m}$ for the hidden layer, and the weight matrix $W^{(2)}\in\mathbb{R}^{m\times d}$ for an output layer. With the popular nonlinear activation ReLU $\sigma$, the GN maps any input $x\in\mathbb{R}^{n}$ to the output: $F(x)=W^{(2)}\sigma(W^{(1)}x)$. For the brevity, we denote by $\alpha_{i}^{\top}$ the $i$-th column of $W^{(1)}$, by $\beta_{j}$ the $j$-th row of $W^{(2)}$ and $W^{(1)}$, and collect these parameters (with $m$ neurons) into a set $\Theta_{m}$. In this formulation, the output of a GN can be formulated as follows:
@@
General equivariant networks (GENs).
GENs represent a subset of GNs that constrain the output function $F$ to be equivalent with respect to a given group representation. A group representation for a group $G$ is formally defined as a homomorphism $\rho:G\to GL(m)$, where $GL(m)$ is the group of all invertible matrices on $\mathbb{R}^{m\times m}$. A function $F$ is considered equivariant if, for given input and output representations $\rho$ and $\phi$ respectively, it adheres to the condition $F\circ\rho_{g}=\phi_{g}\circ F$ for all group elements $g\in G$. A GN whose function $F$ satisfies this property is categorised as a GEN. Specifically, when $\phi=\text{id}$, the function $F$ is called an invariant function and the networks are called invariant networks. In this paper, we assume that group $G$ is finite.
Layer-wise equivariant networks (LENs).
LENs represent a subset of GNs that constrain each layer to be equivalent with respect to a given group representation. Specifically, given the group representation $\psi$ in the hidden layer, LENs require that for all $g\in G$,
@@
As each layer is constrained to be equivariant to maintain the symmetry in the data, the whole network is equivariant. Specifically, for every group element $g\in G$, we have
$F\circ\rho_{g}$ $=W^{(2)}\circ\sigma\circ W^{(1)}\circ\rho_{g}=W^{(2)}\circ\sigma\circ\psi_{g}\circ W^{(1)},$
$=W^{(2)}\circ\psi_{g}\circ\sigma\circ W^{(1)}=\phi_{g}\circ W^{(2)}\circ\sigma\circ W^{(1)}=\phi_{g}\circ F.$
In practice, constructing LENs is a common approach to achieving GENs (he2021efficient).
Expressive power.
The expressive power, of GNs, GENs, and LENs, is defined as the minimum expected loss for approximating a target function $s$. Natually, a smaller expected loss means a better expressive power. Since GENs are in a subset of GNs, and LENs have intuitively more constraints, we have the following straightforward results: for any target function $s$, any input distribution $P\in\Delta\mathbb{R}^{n}$, and any number of neurons $m$,
@@
where $\theta$ represents the parameter of a two-layer ReLU network ($\theta=(W^{(2)},W^{(1)})$), and $F_{\theta}$ is the forward function given parameter $\theta$, and $\Theta_{m}$ is the parameter set of $m$-hidden-layer networks. We reuse the notations GNs (also for GENs and LENs) to denote the respective parameter sets that makes $F_{\theta}$ a GN (also for GEN and LEN). A study on whether equivariance limits the expressive power of two-layer network is thus to study whether there exists some equivariant function $s$ and $m$ such that
$\inf_{\theta\in\Theta_{m}\cap\text{GNs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]$ $<\inf_{\theta\in\Theta_{m}\cap\text{GENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}],\text{ and}$
$\inf_{\theta\in\Theta_{m}\cap\text{GENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}]$ $<\inf_{\theta\in\Theta_{m}\cap\text{LENs}}\mathbb{E}_{x\sim P}[\|F_{\theta}(x)-s(x)\|_{2}^{2}].$
The first inequality studies whether being equivariant is necessary for being the optimal approximator if the target function is equivariant. The second is on whether enforcing layer-wise equivariance is the optimal way to incorporate equivariance. We also note that it is reasonable to assume the target function $s$ to be equivariant, because equivariant networks are designed for equivariant tasks.
4 Boundary Hyperplanes and Channel Vectors
To establish a clear foundation for our analysis of expressive power, we discuss two tools, symmetry boundary hyperplanes and symmetry channel vectors, in this section.
4.1 GENs Imply Symmetric Boundary Hyperplanes
Naturally, ReLU networks are also piece-wise linear, given ReLU functions are piece-wise linear:
@@ 4.1 GENs Imply Symmetric Boundary Hyperplanes
where $\mathcal{F}(x)=\sum_{i:\left\langle\alpha_{i},x\right\rangle\geq 0}^{m}\alpha_{i}\beta_{i}^{\top}$ is the feature function of $F$ and is piece-wise constant.
@@
It is not necessary to define the feature function $\mathcal{F}$ on the boundary hyperplanes. For any given function $F$, we can define it only on the set $S=\{x:F\text{ is linear in }B(x,r),\exists r>0\}$, where $B(x,r)=\{x+z:\|z\|<r\}$ is the open ball centering at $x$ with radius $r$. The value of $\mathcal{F}(x)$ is determined uniquely by the Riesz representation theorem. The complementary set $S^{c}$ is the union of all boundary hyperplanes, whose intersection with the hyperplanes $M\pm\delta v$. For all $x\in S$ that satisfies $\langle x,\alpha_{i}\rangle\not=0$ for all $i\in[m]$, $\mathcal{F}(x)$ must be $\sum_{i:\langle x,\alpha_{i}\rangle>0}\beta_{i}\alpha_{i}^{T}$. Therefore, if the feature function changes when crossing through the same hyperplane, the active set $\{i:\langle x,\alpha_{i}\rangle>0\}$ must change and the increment of the feature function is $\sum_{i:\langle\alpha_{i},x\rangle=0}\beta_{i}\alpha_{i}^{T}$, which is then summarized as Theorem 4.2. This theorem can be viewed as an equivalent definition by the channel vectors.
@@
Boundary hyperplane provides a powerful analytical tool for understanding the internal mechanisms of GENs. We prove that for a GEN, the set of its boundary hyperplanes are symmetric with respect to all group representations. Specifically, for any given boundary hyperplane $M$ and all fixed group representation $\rho$, applying the transformation $\rho$ to the boundary hyperplane (denoted as $\rho M=\{\rho x:x\in M\}$) also gets a boundary hyperplane because
@@
4.2 LENs further Imply Symmetric Channel Vectors
LENs represent a more constrained subset of GENs, enforcing symmetry at the layer level rather than across the network as a whole. Although the symmetry property for boundary hyperplane is a fundamental property of all GENs, the stricter conditions imposed by the LEN structure provide deeper insights. At least, as mentioned in Section 3, LEN needs one admitted group representation and two weight matrices in the corresponding intertwining space.
We first consider the property of admitted group representation, i.e., $\psi\circ\sigma=\sigma\circ\psi$. This commutation property is not trivial and places a significant restriction on the form that the group representation $\psi$ can take. Specifically, for $\sigma=\operatorname{ReLU}$, the following lemma provides a complete characterization of all such admitted representations.